OpenStudy (anonymous):

h(x)=-1/2x-2 if x not equal 2 h(x)=-1 if x equal 2 find the following h(-3)=

OpenStudy (anonymous):

@gelidium

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@MekaylaCox @juliannescriv @xapproachesinfinity

OpenStudy (anonymous):

-3 isnt equal to 2, so you would use h(x)=-1/2x-2

OpenStudy (anonymous):

do I add -3 into that equation

OpenStudy (anonymous):

h(-3)=-1/2(-3)-2

OpenStudy (anonymous):

i got 5/6

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

no, -1/2

OpenStudy (anonymous):

-1/2 * -3 is 3/2

OpenStudy (anonymous):

oh ok thank. how did you get -1/2

OpenStudy (anonymous):

(3/2)-2 = -1/2

OpenStudy (anonymous):